[Brahms-l] Shift report 20040131 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 31 2004 - 23:45:12 EST
------------------------------------------------------------
Date: Sat Jan 31 23:45:12 2004, Shift: 16:00-24:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 47 - 47

Shift Summary:
------------------------------------------------------------
1. Beam for 5 hours.

2. Data runs: 9988, 9989, 9990, 9991, 9992.
  Pedestal runs: 9993 (TPC), 9994 (Non-TPC).

  I finished the B1219(6 deg) setting for FS and I taked 
another ~770k trig6 events for the next setting (D1@B861).
MRS setting was D5@B2255(40deg). I taked 530k trig3 events.

3. No detector problems.
  
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 31 23:46:06 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 31 2004 - 23:46:24 EST